### v1

@@ Abstract
In recent years, generative artificial intelligence models, represented by Large Language Models (LLMs) and Diffusion Models (DMs), have revolutionized content production methods. These artificial intelligence-generated content (AIGC) have become deeply embedded in various aspects of daily life and work, spanning texts, images, videos, and audio. The authenticity of AI-generated content is progressively enhancing, approaching human-level creative standards. However, these technologies have also led to the emergence of Fake Artificial Intelligence Generated Content (FAIGC), posing new challenges in distinguishing genuine information. It is crucial to recognize that AIGC technology is akin to a double-edged sword; its potent generative capabilities, while beneficial, also pose risks for the creation and dissemination of FAIGC. In this survey, We propose a new taxonomy that provides a more comprehensive breakdown of the space of FAIGC methods today. Next, we explore the modalities and generative technologies of FAIGC, categorized under AI-generated disinformation and AI-generated misinformation. From various perspectives, we then introduce FAIGC detection methods, including Deceptive FAIGC Detection, Deepfake Detection, and Hallucination-based FAIGC Detection. Finally, we discuss outstanding challenges and promising areas for future research.
@@ 1 Introduction
  • Innovative and Structured Taxonomy for FAIGC. Our classification and discussion of FAIGC encompass three distinct dimensions: the underlying intent of FAIGC creation, the modal types, and the generative technologies and methodologies. This multi-dimensional approach facilitates a thorough and novel understanding of FAIGC.
@@
  • Future Research Directions in FAIGC. We explore several promising avenues for future research in FAIGC detection. These include diverse areas such as multimodal FAIGC Detection, Zero-shot FAIGC Detection, and interpretable FAIGC Detection.
@@ 6 FAIGC Detection Benchmark
ADD 2023 LR 112003 Chinese F1-score,EER,A-sentence
ADD 2023 AR 110290 Chinese F1-score,EER,A-sentence
@@ 6.1 Deceptive FAIGC Detection
The Deceptive FAIGC detection benchmark focuses on assessing the capacity of detection methods to correctly identify deceptive content within AI-generated responses. It requires detection methods to discern fabricated facts, consistency errors, and similar content. This will test the detection methods’ capabilities in factual retrieval, semantic understanding, and more. In this section, we will explore the corresponding benchmarks from both text and multimodal perspectives.
@@ 6.1.1 Text Modality
News-style GROVER-generated dataset [286]. This benchmark consists of two parts: one part is "real news," which includes 10,000 human-written news articles dated before March 2019, and the other part consists of 10,000 news articles generated by GROVER-Mega. Combining these two parts yields a training set of 10,000 samples, with 2,000 for validation and 8,000 for testing. For the evaluation process, this benchmark employs the Accuracy metric.
Webtext-style GPT2-generated dataset [175]. This benchmark is comprised of two parts: one consists of human-written instances collected from WebText, and the other part consists of Machine-generated instances produced by GPT-2 XL-1542M. This includes a training set of 500,000 samples, a validation set of 10,000, and a test set of 10,000. Common metrics such as F1 and Accuracy can be used for evaluating this benchmark.
Fake News dataset [218]. This benchmark contains 387,000 various news articles collected from the internet, along with news texts generated by OpenAI’s GPT-2 language model. This benchmark can be evaluated using common metrics like F1 and Accuracy.
LLMFake [28]. This benchmark is themed around fake content generated by LLMs. For the construction process, three disinformation generation methods are utilized: Hallucinated News, Arbitrary, and Controllable Generation. Hallucinated and Arbitrary methods prompt models like ChatGPT to produce fake news. Controllable Generation employs prompts on models, such as Llama and Vicuna, to create disinformation from nonfactual content in Politifact, Gossipcop, and CoAID datasets using techniques like Paraphrasing, Rewriting, Open-ended, and Information Manipulation Generation, with the last one solely using ChatGPT and Politifact. This results in approximately 5,400 samples. For the evaluation process, the Accuracy metric is used to assess the detection methods’ ability to identify disinformation.
UHGEval [135]. This benchmark focuses on the recognition of hallucinated content in the news domain. During the construction process, open-source models are used to generate continuations based on given news headlines and openings (sourced from Xinhua News Agency), which typically contain hallucinated content. Subsequently, GPT-3.5 is used to extract keywords included in the response text, and GPT-4 is used to determine the correctness of these keywords. After undergoing a manual review process, this benchmark innovatively proposed the concept of keyword correctness and ultimately produced approximately 5,000 samples. For the evaluation process, in addition to conventional metrics, this benchmark innovatively proposes a keyword precision metric to assess the method’s ability to detect hallucinations at the keyword level.
@@ 6.1.2 Multimdoal
VOA-KG2txt [66]. This benchmark includes a substantial dataset of multimedia fake news, characterized by its fine-grained annotations, which enable the detection of specific elements within the news. The construction process is divided into two methods. The first method involves generating fake news from a KG of a real article, manipulating mid-frequency entities through swapping, adding relations, or replacing subgraphs, and using BART-large for text generation. The second method constructs a KG and uses an Abstract Meaning Representation (AMR) parser for graph manipulation via role switching or predicate negation. This process results in a dataset comprising 15k real news and 15k machine-generated fake news. For the evaluation process, the benchmark employs the Accuracy metric to assess the performance of detection methods at the document level, and the F-score metric to evaluate detection methods at the fine-grained Knowledge Entity (KE) level.
NeuralNews [230]. This benchmark contains a vast collection of both human and machine-generated articles, along with images and captions. Machine-generated news articles are created by re-writing human articles using models such as GROVER. The data in this benchmark is categorized into four forms: (1) Real Articles with Real Captions, (2) Real Articles with Generated Captions, (3) Generated Articles with Real Captions, and (4) Generated Articles with Generated Captions, each with 32K samples. For the evaluation process, this benchmark uses the Accuracy metric to evaluate the performance of detection methods.
@@ 6.2.1 LLMs Hallucination
FactCHD [35]. This benchmark focuses on identifying fact-conflicting hallucinated content in LLMs’responses. To construct the benchmark, manually designed queries are used to generate query-response pairs with open-source LLMs (ChatGLM, Vicuna, Alpca), which are annotated as non-factual and factual to serve as examples. Prompts are then designed for ChatGPT to learn from these examples and generate similar query-response pairs, with data selection based on Semantic Diversity. Further prompts lead ChatGPT to generate explanations for these pairs, culminating in a dataset of 6,940 samples after manual curation. For the evaluation process, this benchmark utilizes the F1 score to assess the detection methods’ ability to identify hallucinations and a custom metric similar to the n-gram evaluation metric to evaluate the ability to generate corresponding reasonable explanations.
@@
PHD [273]. This benchmark focuses on the identification of hallucinated content at the passage level. It is constructed by extracting entities from a Wikipedia dump, dividing them into three categories based on the number of Google Search results (PHD-Low, PHD-Medium, PHD-High), and sampling 100 entities from each category. Prompts guide ChatGPT to generate Wikipedia-style articles, which are manually labeled as factual, non-factual, or unverifiable. For the evaluation process, this benchmark utilizes Precision, Recall, and F1 metrics to assess the detection method’s ability to identify passage-level hallucinations.
@@
AutoHall [21]. This benchmark is concentrated on the detection of fact-conflicting hallucinations. The construction process involves creating prompts for an LLM to generate references for claims labeled as True or False. The LLM then assesses the claim’s correctness based on the provided claim and reference. If the assessment contradicts the original label, the previously generated reference is considered to contain hallucination. Data for this benchmark are sourced from three known datasets: Climate-fever, Pubhealth, and WICE, resulting in a total of 907, 1009, and 928 samples, respectively. For the evaluation process, this benchmark utilizes the Accuracy and F1 score metrics to assess the detection method’s capability in identifying hallucinations.
@@
HADES [145]. This benchmark addresses token-level hallucination detection. It is constructed by sampling text segments from the WIKI-40B dataset, perturbing them with BERT while adhering to fluency, grammatical correctness, and diversity principles. The perturbed texts are manually annotated as hallucination or not, yielding a dataset of 10,954 samples. For the evaluation process, this benchmark employs Precision, Recall, and F1 score metrics to evaluate the detection method’s ability to identify token-level hallucinated content.
@@
RAGTruth [261]. This benchmark is focused on hallucination detection during the RAG process. For the construction process, responses are generated for the RAG task using six models, subsequently annotated manually as Yes (hallucination) or No (not hallucination). The RAG tasks are divided into three categories: Summarization, Question Answering, and Data-to-Text generation, with data sourced from CNN/Daily Mail, MS MACRO, and Yelp Open Dataset, respectively, amounting to a total of 17,838 samples. For the evaluation process, this benchmark employs Precision, Recall, and F1 score metrics to evaluate the detection method’s ability to identify token-level hallucinations.
@@
DelucionQA [203]. This benchmark focuses on hallucination detection in domain-specific QA tasks during the RAG process. It involves using an LLM to generate questions about the Jeep 2023 Gladiator dataset, followed by manual filtering to remove inappropriate questions. A retrieval algorithm provides evidence for these questions, which the LLM uses to generate responses. These responses are manually annotated, resulting in 2,038 samples. For the evaluation process, this benchmark employs the F1 score metric to assess the detection method’s capability to identify hallucinations.
@@
HaluEval [126]. This benchmark focuses on the identification of hallucinated content within NLG tasks. The construction method is divided into automated generation and human annotation. For the automated generation process, using a "sampling-then-filtering" approach prompts ChatGPT to generate hallucinated responses, and then refine them to plausible answers. This generates 10,000 samples each for question answering (HotpotQA), text summarization (CNN/Daily Mail), and knowledge-grounded dialogue (OpenDialKG). For human annotation, queries are sampled from the 52K instruction tuning dataset from Alpaca, and ChatGPT is used to generate corresponding responses. These are manually labeled as unverifiable, non-factual, and irrelevant, resulting in a dataset of 5,000 samples. For the evaluation process, this benchmark employs the Accuracy metric to assess the detection method’s capability to identify hallucinations.
@@
SelfCheckGPT-Wikibio [155]. This benchmark focuses on the detection of fact-conflicting hallucinated content. The construction process involves sampling a set of individuals/concepts from the WikiBio dataset and creating prompts that instruct ChatGPT to generate Wikipedia-style articles. These articles are subsequently manually annotated, with each sentence being labeled as either Major Inaccurate, Minor Inaccurate, or Accurate. This process results in a dataset comprising 1,908 samples. For the evaluation process, the benchmark employs the AUROC metric to assess the detection method’s ability to identify hallucinations.
@@
FAVABENCH [160]. This benchmark concentrates on the detection of hallucinated content within fine-grained factuality tasks. The construction process involves collecting authentic responses from mainstream models (such as ChatGPT, Llama2-Chat 7B, and Llama2-Chat 70B) across multiple knowledge-intensive tasks. These responses are manually annotated with fine granularity (such as entities, relations, etc.), resulting in a dataset of approximately 1,000 samples. For the evaluation process, the benchmark employs the F1 score to assess the detection method’s ability to identify hallucinations.
@@ 6.2.2 Multimodal
MHaluBench [36]. The benchmark comprises two types of data: text-to-image and image-to-text. In the construction phase, for the former, the dataset includes 200 data points from Image Captioning (IC) tasks and 200 from Visual Question Answering (VQA) tasks. For the latter, it contains 220 data points from Image-to-Text (I2T) tasks. The data is generated by state-of-the-art MLLMs such as LLaVA, MiniGPT-4, and DALLE-3. This benchmark introduces a multi-granular evaluation approach, where the input text and output text are segmented into multiple segments and claims to achieve a more accurate assessment. The annotation is done manually. For the evaluation process, this benchmark employs Precision, Recall, F1, and Accuracy metrics to evaluate the detection method’s ability to identify hallucinations.
HalDetect [76]. This benchmark focuses on VQA tasks. The construction process involves sampling 4,000 images from the COCO dataset and generating descriptive responses to those images using the InstructBLIP model. For each image, four responses are sampled using nucleus sampling, resulting in a total of 16,000 samples. The dataset is manually annotated with a fine-grained approach, which entails segmenting each sentence into various fragments and labeling them. The annotations are categorized into four types: Accurate, Inaccurate, Analysis, and Unsure. ’Analysis’ represents some form of reasoning or explanatory process, which is typically irrelevant to the provided image. For the evaluation process, this benchmark utilizes F1, and Accuracy metrics to evaluate the detection method’s ability to identify hallucinations.
@@ 6.3.1 Visual Deepfake Detect
CDDB [124]. This benchmark extends the stationary Deepfake detection scenario to a dynamic one, in which a stream of likely heterogeneous Deepfake appears over time rather than simultaneously. In this setting, previously encountered Deepfake may become invisible due to certain reasons. This poses a greater challenge to current methods that are more inclined toward static detection. The construction process involves using three types of models: 1) GAN models, 2) non-GAN models, and 3) unknown models to synthesize images from a series of real sources, ultimately yielding approximately 842k samples. For the evaluation process, this benchmark employs AA, AF, AA-M, and mAP metrics to assess the performance of detection methods.
@@
FakeAVCeleb [104]. This research introduces an innovative audiovisual Deepfake dataset, termed FakeAVCeleb. Comprising not only Deepfake videos, this dataset distinctively features associated synthetic audio that is lip-sync synchronized with the forged visual content. Data generation was conducted utilizing prevalent Deepfake production methods, with a particular focus on selecting authentic YouTube videos of celebrities representing diverse racial backgrounds, aiming to cultivate a more sophisticated and realistic multimodal dataset. Additionally, the study executed a series of experiments to thoroughly evaluate this Deepfake dataset. These investigations shed light on both the challenges and potential applications of this multimodal audiovisual Deepfake dataset in the realm of detection.
@@
Deepfake MNIST+ [85]. This research introduces a novel facial animation dataset named Deepfake MNIST+1, designed to tackle the pivotal and challenging role of facial animations in Deepfake detection. Generated by state-of-the-art image animation generators, this dataset comprises 10,000 facial animation videos, encompassing a diverse range of ten distinct facial expressions or actions. Additionally, the study presents a baseline detection methodology, followed by an in-depth analysis of this approach. Furthermore, it examines the characteristics of the newly proposed dataset, highlighting the complexities and critical importance of identifying Deepfake videos amidst varying facial expressions and different levels of compression quality.

### v2

@@ Abstract
In recent years, generative artificial intelligence models, represented by Large Language Models (LLMs) and Diffusion Models (DMs), have revolutionized content production methods. These artificial intelligence-generated content (AIGC) have become deeply embedded in various aspects of daily life and work. However, these technologies have also led to the emergence of Fake Artificial Intelligence Generated Content (FAIGC), posing new challenges in distinguishing genuine information. It is crucial to recognize that AIGC technology is akin to a double-edged sword; its potent generative capabilities, while beneficial, also pose risks for the creation and dissemination of FAIGC. In this survey, We propose a new taxonomy that provides a more comprehensive breakdown of the space of FAIGC methods today. Next, we explore the modalities and generative technologies of FAIGC. We introduce FAIGC detection methods and summarize the related benchmark from various perspectives. Finally, we discuss outstanding challenges and promising areas for future research.
@@ 1 Introduction
  • Innovative and Structured Taxonomy for FAIGC. Our classification and discussion of FAIGC encompass three distinct dimensions: the intent behind FAIGC, the modalities and generative technologies of FAIGC, and the creation method of FAIGC. This multi-dimensional approach facilitates a thorough and novel understanding of FAIGC.
@@
  • Future Research Directions in FAIGC. We explore several promising avenues for future research in FAIGC detection. These include diverse areas such as multimodal FAIGC Detection, Zero-shot FAIGC Detection, and interpretable FAIGC Detection and so on. It provides valuable ideas for future researchers in this field.
@@ 6 FAIGC Detection Benchmark
ADD 2023[276] LR 112003 Chinese F1-score,EER,A-sentence
ADD 2023[276] AR 110290 Chinese F1-score,EER,A-sentence
@@ 6.1 Deceptive FAIGC Detection
The Deceptive FAIGC detection benchmark focuses on assessing the capacity of detection methods to identify deceptive content within AI-generated responses correctly. It requires detection methods to discern fabricated facts, consistency errors, and similar content. This will test the detection methods’ capabilities in factual retrieval, semantic understanding, and more. In this section, we will explore the corresponding benchmarks from both text and multimodal perspectives.
@@ 6.1.1 Text Modality
News-style GROVER-generated dataset [286]. This benchmark consists of two parts: one part is "real news," which includes 10,000 human-written news articles dated before March 2019, and the other part consists of 10,000 news articles generated by GROVER-Mega. Combining these two parts yields a training set of 10,000 samples, with 2,000 for validation and 8,000 for testing.
Webtext-style GPT2-generated dataset [175]. This benchmark is comprised of two parts: one consists of human-written instances collected from WebText, and the other part consists of Machine-generated instances produced by GPT-2 XL-1542M. This includes a training set of 500,000 samples, a validation set of 10,000, and a test set of 10,000.
Fake News dataset [218]. This benchmark contains 387,000 various news articles collected from the internet, along with news texts generated by OpenAI’s GPT-2 language model.
LLMFake [28]. This benchmark is themed around fake content generated by LLMs. For the construction process, three disinformation generation methods are utilized: Hallucinated News, Arbitrary, and Controllable Generation. Hallucinated and Arbitrary methods prompt models like ChatGPT to produce fake news. Controllable Generation employs prompts on models, such as Llama and Vicuna, to create disinformation from nonfactual content in Politifact, Gossipcop, and CoAID datasets using techniques like Paraphrasing, Rewriting, Open-ended, and Information Manipulation Generation, with the last one solely using ChatGPT and Politifact. This results in approximately 5,400 samples.
UHGEval [135]. This benchmark focuses on the recognition of hallucinated content in the news domain. During the construction process, open-source models are used to generate continuations based on given news headlines and openings, which typically contain hallucinated content. Subsequently, GPT-3.5 is used to extract keywords from the response, GPT-4 is used to determine the correctness of these keywords, and then undergoes a manual review process. Ultimately resulting in approximately 5,000 samples.
@@ 6.1.2 Multimdoal
VOA-KG2txt [66]. This benchmark includes a substantial dataset of multimedia fake news, characterized by its fine-grained annotations, which enable the detection of specific elements within the news. The construction process is divided into two methods. The first method involves generating fake news from a KG of a real article, manipulating mid-frequency entities through swapping, adding relations, or replacing subgraphs, and using BART-large for text generation. The second method constructs a KG and uses an Abstract Meaning Representation (AMR) parser for graph manipulation via role switching or predicate negation. This process results in a dataset comprising 15k real news and 15k machine-generated fake news.
NeuralNews [230]. This benchmark contains a vast collection of both human and machine-generated articles, along with images and captions. Machine-generated news articles are created by re-writing human articles using models such as GROVER. The data in this benchmark is categorized into four forms: (1) Real Articles with Real Captions, (2) Real Articles with Generated Captions, (3) Generated Articles with Real Captions, and (4) Generated Articles with Generated Captions, each with 32K samples.
@@ 6.2.1 LLMs Hallucination
FactCHD [35]. This benchmark focuses on identifying fact-conflicting hallucinated content in LLMs’responses. To construct the benchmark, manually designed queries are used to generate query-response pairs with open-source LLMs, which are annotated as non-factual and factual to serve as examples. Prompts are then designed for ChatGPT to learn from these examples and generate similar query-response pairs, with data selection based on Semantic Diversity. Further prompts lead ChatGPT to generate explanations for these pairs, culminating in a dataset of 6,940 samples after manual curation.
@@
PHD [273]. This benchmark focuses on the identification of hallucinated content at the passage level. It is constructed by extracting entities from a Wikipedia dump, dividing them into three categories based on the number of Google Search results (PHD-Low, PHD-Medium, PHD-High), and sampling 100 entities from each category. Prompts guide ChatGPT to generate Wikipedia-style articles, which are manually labeled as factual, non-factual, or unverifiable.
@@
AutoHall [21]. This benchmark is concentrated on the detection of fact-conflicting hallucinations. The construction process involves creating prompts for an LLM to generate references for claims labeled as True or False. The LLM then assesses the claim’s correctness based on the provided claim and reference. If the assessment contradicts the original label, the previously generated reference is considered to contain hallucination. Data for this benchmark are sourced from three known datasets: Climate-fever, Pubhealth, and WICE, resulting in a total of 907, 1009, and 928 samples, respectively.
@@
HADES [145]. This benchmark addresses token-level hallucination detection. It is constructed by sampling text segments from the WIKI-40B dataset, perturbing them with BERT while adhering to fluency, grammatical correctness, and diversity principles. The perturbed texts are manually annotated as hallucination or not, yielding a dataset of 10,954 samples.
@@
RAGTruth [261]. This benchmark is focused on hallucination detection during the RAG process. For the construction process, responses are generated for the RAG task using six models, subsequently annotated manually as Yes (hallucination) or No (not hallucination). The RAG tasks are divided into three categories: Summarization, Question Answering, and Data-to-Text generation, with data sourced from CNN/Daily Mail, MS MACRO, and Yelp Open Dataset, respectively, amounting to a total of 17,838 samples.
@@
DelucionQA [203]. This benchmark focuses on hallucination detection in domain-specific QA tasks during the RAG process. It involves using an LLM to generate questions about the Jeep 2023 Gladiator dataset, followed by manual filtering to remove inappropriate questions. A retrieval algorithm provides evidence for these questions, which the LLM uses to generate responses. These responses are manually annotated, resulting in 2,038 samples.
@@
HaluEval [126]. This benchmark focuses on the identification of hallucinated content within NLG tasks. The construction method is divided into automated generation and human annotation. For the former, using a "sampling-then-filtering" approach prompts ChatGPT to generate hallucinated responses, and then refine them to plausible answers. This generates 10,000 samples each for question answering, text summarization, and knowledge-grounded dialogue. For human annotation, queries are sampled from the 52K instruction tuning dataset from Alpaca, and ChatGPT is used to generate corresponding responses. These are manually labeled as unverifiable, non-factual, and irrelevant, resulting in a dataset of 5,000 samples.
@@
SelfCheckGPT-Wikibio [155]. This benchmark focuses on the detection of fact-conflicting hallucinated content. The construction process involves sampling a set of individuals/concepts from the WikiBio dataset and creating prompts that instruct ChatGPT to generate Wikipedia-style articles. These articles are subsequently manually annotated, with each sentence being labeled as either Major Inaccurate, Minor Inaccurate, or Accurate. This process results in a dataset comprising 1,908 samples.
@@
FAVABENCH [160]. This benchmark concentrates on the detection of hallucinated content within fine-grained factuality tasks. The construction process involves collecting authentic responses from mainstream LLMs across multiple knowledge-intensive tasks. These responses are manually annotated with fine granularity (such as entities, relations, etc.), resulting in a dataset of approximately 1,000 samples.
@@ 6.2.2 Multimodal
MHaluBench [36]. The benchmark comprises two types of data: text-to-image and image-to-text. In the construction phase, for the former, the dataset includes 200 data points from Image Captioning (IC) tasks and 200 from Visual Question Answering (VQA) tasks. For the latter, it contains 220 data points from Image-to-Text (I2T) tasks. The data is generated by state-of-the-art MLLMs such as LLaVA, MiniGPT-4, and DALLE-3. This benchmark introduces a multi-granular evaluation approach, where the input text and output text are segmented into multiple segments and claims to achieve a more accurate assessment. The annotation is done manually.
HalDetect [76]. This benchmark focuses on VQA tasks. The construction process involves sampling 4,000 images from the COCO dataset and generating descriptive responses to those images using the InstructBLIP model. For each image, four responses are sampled using nucleus sampling, resulting in a total of 16,000 samples. The dataset is manually annotated with a fine-grained approach, which entails segmenting each sentence into various fragments and labeling them. The annotations are categorized into four types: Accurate, Inaccurate, Analysis, and Unsure. ’Analysis’ represents some form of reasoning or explanatory process, which is typically irrelevant to the provided image.
@@ 6.3.1 Visual Deepfake Detect
CDDB [124]. This benchmark extends the stationary Deepfake detection scenario to a dynamic one, in which a stream of likely heterogeneous Deepfake appears over time rather than simultaneously. In this setting, previously encountered Deepfake may become invisible due to certain reasons. This poses a greater challenge to current methods that are more inclined toward static detection. The construction process involves using three types of models: 1) GAN models, 2) non-GAN models, and 3) unknown models to synthesize images from a series of real sources, ultimately yielding approximately 842k samples.
@@
FakeAVCeleb [104]. This research introduces FakeAVCeleb, an innovative audiovisual Deepfake dataset. Unlike others, it includes synthetic audio synchronized with the forged visuals, created using prevalent Deepfake methods. The dataset features diverse celebrity videos to enhance realism. The study conducts thorough experiments to evaluate it, revealing challenges and potential detection applications.
@@
Deepfake MNIST+ [85]. This research introduces Deepfake MNIST+1, a novel dataset containing 10,000 facial animation videos. It covers ten diverse facial expressions/actions and serves to aid in Deepfake detection. The study includes a baseline detection method and an analysis of its effectiveness. Furthermore, it examines the characteristics of the newly proposed dataset, highlighting the complexities and critical importance of identifying Deepfake videos amidst varying facial expressions and different levels of compression quality.